# Gravity as a deformed topological gauge theory

### v6

@@ 2.3 Associated action
These dependencies on $G,\gamma$, $\Lambda_{0}$ and $r$ are interesting since a quantum version of the theory may yield effects sensitive to the bare cosmological constant $\Lambda_{0}$ in particular. We believe that probing the bare cosmological constant in this way could be highly beneficial in constraining the vacuum contribution to the cosmological constant $\Lambda_{\text{vac}}$ since the two cases $\Lambda_{\text{vac}}=0$ and $\Lambda=\Lambda_{\text{vac}}=8\pi G\rho_{vac}$ bring about huge differences in the numerical value of $\Lambda_{0}$.
One can also see that the coupling constant of the Euler density $-\dfrac{3\alpha_{1}}{4\Lambda_{0}}R^{ab}\wedge R^{cd}\varepsilon_{abcd}$ obtained via the equation of motion (9) for the tetrad $\beta$ matches perfectly with what has been described in [12, section III] such that the gauge action $S_{G}$ (15) is invariant under the "duality symmetry" consisting of the exchange $R^{ab}\leftrightarrow\dfrac{\Lambda_{0}}{3}\beta^{a}\wedge\beta^{b}$. It is worthwhile to notice that since by construction $S_{G}$ is quadratic in the curvature, $S_{G}$ turns out to be invariant under the (more general) duality symmetry $\bar{\Omega}\leftrightarrow-\bar{\Omega}$ which splits as:
@@
In addition, a direct comparison with the two perturbative expansion parameters $\alpha$,$\beta$ given in [18] yields:
@@
As for $k$, the case of a dynamical $\ell$ will be studied in a forthcoming paper.
[…]
@@
The Pontrjagin density and other terms described in (15) correspond to the Chern-Simons modified gravity [25] terms with a constant scalar field $\vartheta=-\big{(}\dfrac{2Gr}{\pi}+\dfrac{3}{\Lambda_{0}\gamma}\big{)}\dfrac{\alpha_{1}}{2}$ and potential term for $\vartheta$, ${\displaystyle S_{\vartheta}=\int_{\mathscr{M}}V(\vartheta)}$, with:
@@
The Chern-Simons term also occurs in Loop Quantum Gravity (LQG) and String theory as mentionned in [25].
@@
3 Invariant polynomials and action for the Möbius group $G=SO(4,2)/\{\pm I\}$
@@
The generators of the above graded algebra are the same as the ones described previously in (2.1) except for $\mathfrak{g}_{0}\simeq\mathfrak{co}(3,1)$ which is generated by the $\{J_{ab}\}$’s and $\mathds{1}_{4}$. An element of $\mathfrak{co}(3,1)\simeq\mathfrak{g}_{0}$ is $\gamma_{\mathfrak{co}}=\frac{1}{2}\gamma_{\mathfrak{so}}^{ab}J_{ab}-z\mathds{1}_{4}$ with $\gamma_{\mathfrak{so}}\in\mathfrak{so}(3,,1)$ and $z\in\mathbb{R}$.
3.1 The Möbius Lie algebra as a symmetric Lie algebra
Let us consider $H=CO(3,1)$ as structure group for the Cartan geometry to be discussed in this part. This implies a geometry that differs from the whole isotropic Möbius group as introduced in [11, Lemma 1.8,p.269]). It rather corresponds to $SO(3,1)\times\mathbb{R}^{+}$ (Lorentz×$\times$×Weyl) the isotropic part of the Weyl model which can be embedded into the Möbius model, see [11, middle p.277].
@@ 3.1 The Möbius Lie algebra as a symmetric Lie algebra
of the Möbius group is reductive with respect to $H$ and is also a symmetric Lie algebra $\mathfrak{g}=\mathfrak{co}(3,1)\oplus\mathfrak{m}$ where $\mathfrak{m}=\mathfrak{g}_{1}\oplus\mathfrak{g}_{-1}\simeq{\mathbb{R}^{3,1}}^{*}\oplus\mathbb{R}^{3,1}$. The commutation relations (see (LABEL:eq:gradedLa) and [14, Example 5.2 and Chapter XI]) are given by
$[\mathfrak{g}_{0},\mathfrak{g}_{0}]\subset\mathfrak{g}_{0},\qquad[\mathfrak{g}_{0},\mathfrak{m}]\subset\mathfrak{m},\qquad[\mathfrak{m},\mathfrak{m}]\subset\mathfrak{g}_{0}$ (2)
@@
for which we adopt the matrix presentation [^14] Accordingly, the Killing form on $\mathfrak{g}=\mathfrak{h}\oplus\mathfrak{m}$ is computed to be $K_{\mathfrak{g}}(\varphi,\varphi^{\prime})=\operatorname{Tr}(ad_{\varphi}\circ ad_{\varphi^{\prime}})=n\operatorname{Tr}(cc^{\prime})+2n\big{(}zz^{\prime}+\eta(b,a^{\prime})+\eta(a,b^{\prime})\big{)}$ where $n$ corresponds to the one coming from $\dim\mathfrak{so}(n-1,1)=n(n-1)/2$.
@@
Let us now consider the mutation [^15] See Definition 2.1 given above. $\mathfrak{g}^{\prime}=\mu(\mathfrak{g})=\mu(\mathfrak{h}\oplus\mathfrak{m})=\mathfrak{h}\oplus\mu(\mathfrak{m})=\mathfrak{h}\oplus\mathfrak{m}^{\prime}$ given by:
@@
Let us choose $\gamma_{1}=\bar{\gamma}_{1}=\dfrac{1}{\ell}$, $\gamma_{2}=\bar{\gamma}_{2}=\dfrac{1}{\ell^{\prime}}$ (where $\ell$ and $\ell^{\prime}$ are non-null real numbers) [^16] In that case, the Killing form on $\mathfrak{g}^{\prime}=\mathfrak{h}\oplus\mathfrak{m}^{\prime}$ can be easily computed by performing the scale transformations $b\mapsto\gamma_{1}b$ and $a\mapsto\gamma_{2}a$ into the expression for $K_{\mathfrak{g}}$, so that $K_{\mathfrak{g}^{\prime}}(\varphi,\varphi^{\prime})=n\operatorname{Tr}(cc^{\prime})+2n\Bigl{(}zz^{\prime}+\gamma_{1}\gamma_{2}\bigl{(}\eta(b,a^{\prime})+\eta(a,b^{\prime})\bigr{)}\Bigr{)}.$
@@ A Cartan connection $\varpi$ with this $G/H$ Cartan geometry can accordingly be decomposed as:
The geometry modeled on the present 888-dimensional quotient $G/H$ describes in general an 888-dimensional manifold $\widetilde{\mathscr{M}}$ invariant under Lorentz transformations (elements in $SO(3,1)$) and dilations (elements of the multiplicative group $\mathbb{R}^{+}$) with soldering form $\varpi_{\mathfrak{m}}$. In the next section, we will see how we can obtain a description of a 444-dimensional submanifold $\mathscr{M}$ of $\widetilde{\mathscr{M}}$ by imposing a condition on the soldering form (the $\mathfrak{m}$-valued part of the Cartan connection).
@@
(see e.g. [26]). For further use, it is worthwhile to notice that the $[\mathfrak{m},\mathfrak{m}]\subset\mathfrak{h}$ contribution to $\bar{\Omega}_{\mathfrak{h}}$ is given by
@@
[…]
3.2 Characteristic poylynomials for $G=SO(4,2)/\{\pm I\}$ and associated action
@@
where $\mathscr{M}_{i}$ is either the 888-dimensional manifold $\widetilde{\mathscr{M}}$ defined by the Cartan geometry $G/H$ or a 444-dimensional submanifold $\mathscr{M}\subset\widetilde{\mathscr{M}}$ defined by the distribution:
Proposition 3.1.
@@
Hence, by the Frobenius theorem [11, Prop.5.3, p.81] the corresponding distribution defines $\mathscr{M}$ as an integrable submanifold of $\widetilde{\mathscr{M}}$. Therefore, there is a subclass of conformal Cartan connections characterizing $\mathscr{M}$; this subclass can be reached through this kind of "gauge fixing".171717 We are very indebted to Thierry Masson for his key input in the construction of this integrable distribution defined by this specific Cartan connection. This ought to resemble a gauge fixing which amounts to freezing the degrees of freedom of the $\mathfrak{g}_{1}$-part to be those of the $\mathfrak{g}_{-1}$-sector. It can be linked with [11, Proof of Proposition 3.1, p.285] where the two tetrads are taken to be linearly dependent each other. To some extent, this selects a mutated Weyl geometry within the mutated Möbius one. One thus gets $\dim\mathscr{M}=11-7=4$. One may also consider manifolds of other dimensions by imposing other constraints that reduce the number of degrees of freedom of the soldering form $\varpi_{\mathfrak{m}}$.
@@
the action being invariant under each of these transformations. Particular transformations satisfying 3 are:
@@
While for $k\neq 0$, a direct comparison between $S^{\prime}$ (4) and $S_{G}$ (3.2) amounts to identifying the 6 coupling constants present in $S^{\prime}$ via the system:
@@
  • Action (3.2) gives back the equations of motion (9)-(10).
  • Adding the term $-\,\dfrac{vk}{2\pi^{2}\ell\ell^{\prime}}\eta_{ab}T^{a}\wedge\lambda\wedge\beta^{b}$ in the action (3.2) leads to a second source of curvature expressed in terms of the spin density of matter, torsion and their variations for the Einstein equations.
3.2.1 Fermionic matter field action
@@
If one considers the usual covariant derivative $D_{\mathfrak{so}}=(\partial_{\mu}+\dfrac{1}{4}A_{ab,\mu}\gamma^{a}\gamma^{b})dx^{\mu}$, the matter action similar to (3) is then given by  [^20] Where we set $\kappa^{2}\zeta=\tfrac{(\ell\ell^{\prime})^{2}}{384k^{2}}$ such that the overall factor is arranged to stick with the Dirac action given in [17, p.197].:
@@
Let $C_{i}$ be the $i$-dimensional Clifford algebra. We can extend the isomorphism 2 between $\mathfrak{so}(n-1,1)$ and $\mathfrak{spin}(n-1,1)$ to the morphism $\Phi_{\varepsilon}:\mathfrak{gl}(n,\mathbb{R})\rightarrow C_{2}\oplus C_{0}$. Let $M=M^{rs}e_{r}\bar{e}_{s}\in\mathfrak{gl(n,\mathbb{R)}}$. We define:
@@ 3.2.2 The total action
In the other case, using the Hodge star operator lets us show that $T^{a}\wedge\beta_{a}=0$ is equivalent to requiring the axial vector part of torsion to vanish ($\tilde{t}^{a}=\dfrac{1}{6}T^{b}_{~\mu\nu}\varepsilon^{a~\mu\nu}_{~b}=0$) as is described in [28, Table 3]. The end result being that the new torsion $T^{a}_{~bc}=Z^{a}_{~bc}+\delta^{a}_{[b}t^{\vphantom{a}}_{c]}$ can be expressed solely in terms of 202020 independent components instead of the usual 242424. Here $Z^{a}_{~bc}$ is a 333-indices tensor and $t_{c}$ is a 444-dimensional vector with respectively 161616 and 444 independent components.
@@
Using the relation $\gamma=\dfrac{2\alpha_{1}}{2\alpha_{2}-\alpha_{5}}$, given in [16, p13], we identify $\gamma=\dfrac{e}{r+2y}$, with $\gamma$ the Barbero-Immirzi parameter.
@@
$G_{kc}+(\dfrac{1}{2\gamma}-\dfrac{\Lambda_{0}Gr}{3\pi})R^{ab}_{~~rs}\varepsilon^{rs}_{~~bk}\eta_{ac}+\Lambda_{0}\eta_{kc}=-\dfrac{1}{2\alpha_{1}}(\tau_{ck}+\dfrac{1}{12}\partial_{\mu}\mathfrak{s}^{ab}_{~~ab\nu}\varepsilon^{\mu\nu}_{~~\;ck}).$ (26)
@@
In this particular case the fact that Torsion is null implies that the Einstein tensor $G_{kc}$ is symmetric (see [29, p3] for more details). When taking into account that $\eta_{kc}$ is symmetric as well, the anti-symmetricity of $\partial_{\mu}\mathfrak{s}^{ab}_{~~ab\nu}\varepsilon^{\mu\nu}_{~~\;ck}$ implies that the energy-momentum tensor $\tau_{ck}$ sports an anti-symmetric part if $\partial_{\mu}\mathfrak{s}^{ab}_{~~ab\nu}\varepsilon^{\mu\nu}_{~~\;ck}\neq 0$. This confirms that we can consider a matter Lagrangian that depends on the spin connection $A$ as we explicitly did.
Separating the energy-momentum tensor $\tau_{ck}$ as $\tau_{ck}=\tau_{M,ck}-\rho_{\text{vac}}\eta_{ck}$ with $-\rho_{\text{vac}}\eta_{ck}$ the part related to the vacuum energy density $\rho_{\text{vac}}\geq 0$ then leads to:
$G_{kc}+(\dfrac{1}{2\gamma}-\dfrac{\Lambda_{0}Gr}{3\pi})R^{ab}_{~~rs}\varepsilon^{rs}_{~~bk}\eta_{ac}+\Lambda\eta_{kc}=-\dfrac{1}{2\alpha_{1}}(\tau_{M,ck}+\dfrac{1}{12}\partial_{\mu}\mathfrak{s}^{ab}_{~~ab\nu}\varepsilon^{\mu\nu}_{~~\;ck})$ (27)
@@
$G_{kc}$ $+(\dfrac{1}{2\gamma}-\dfrac{\Lambda_{0}Gr}{3\pi})R^{ab}_{~~rs}\varepsilon^{rs}_{~~bk}\eta_{ac}+\Lambda_{0}\eta_{kc}$
$=-\dfrac{1}{2\alpha_{1}}\Bigl{(}\tau_{ck}+\dfrac{1}{12}(\eta_{cs}\beta^{s}_{\nu}\partial_{\mu}+T_{c,\mu\nu})\bigl{(}\mathfrak{s}^{ab}_{~~abr}+(\dfrac{1}{2\pi G\gamma}-\dfrac{\Lambda_{0}r}{3\pi^{2}})T^{m}_{~rm}\bigr{)}\varepsilon^{\mu\nu r}_{~~~\,k}\Bigr{)}$ (28)
@@
The secondary source of curvature $-\dfrac{1}{24\alpha_{1}}(\eta_{cs}\beta^{s}_{\nu}\partial_{\mu}+T_{c,\mu\nu})\bigl{(}\mathfrak{s}^{ab}_{~~abr}+(\dfrac{1}{2\pi G\gamma}-\dfrac{\Lambda_{0}r}{3\pi^{2}})T^{m}_{~rm}\bigr{)}\varepsilon^{\mu\nu r}_{~~~\,k}$ in (28) is highly dependent on the value of $\Lambda_{0}$ and $r$. One can for example tune $r=\dfrac{3\pi}{2\Lambda_{0}G\gamma}$ such that it is now dominated by the term $-\dfrac{1}{24\alpha_{1}}(\eta_{cs}\beta^{s}_{\nu}\partial_{\mu}+T_{c,\mu\nu})\mathfrak{s}^{ab}_{~~abr}\varepsilon^{\mu\nu r}_{~~~\,k}$.
@@
Replacing $e=\dfrac{8\pi^{2}\alpha_{1}\ell\ell^{\prime}}{k}=\dfrac{3\pi}{2\Lambda_{0}G}$; $\ell\ell^{\prime}=-\dfrac{3k}{\Lambda_{0}}$ and $y=\dfrac{3\pi}{4\Lambda_{0}G\gamma}-\dfrac{r}{2}$ in the action (3.2), we obtain:
@@
$\qquad\overbrace{-\dfrac{4\Lambda_{0}G\alpha_{1}r}{3\pi}\bigl{(}T^{a}\wedge T_{a}-R^{ab}\wedge\beta_{a}\wedge\beta_{b}\bigr{)}}^{Nieh-Yan}\overbrace{-(\dfrac{r}{8\pi^{2}}+\dfrac{3}{16\pi\Lambda_{0}G\gamma})d\lambda\wedge d\lambda}^{\text{ Dilation kinetic term }}.$ (32)

### v7

@@ II.3 Associated action
These dependencies on $G,\gamma$, $\Lambda_{0}$ and $r$ are interesting since a quantum version of the theory may yield effects sensitive to the bare cosmological constant $\Lambda_{0}$ in particular. We believe that probing the bare cosmological constant in this way could be highly beneficial in constraining the vacuum contribution to the cosmological constant $\Lambda_{\text{vac}}$ since the two cases $\Lambda_{\text{vac}}=0$ and $\Lambda=\Lambda_{\text{vac}}=8\pi G\rho_{\text{vac}}$ bring about huge differences in the numerical value of $\Lambda_{0}$.
One can also see that the coupling constant of the Euler density $\dfrac{3\alpha_{1}}{4\Lambda_{0}}R^{ab}\wedge R^{cd}\varepsilon_{abcd}$ obtained via the equation of motion (36) for the tetrad $\beta$ matches perfectly with what has been described in [13, section III] such that the gauge action $S_{G}$ (42) is invariant under the "duality symmetry" consisting of the exchange $R^{ab}\leftrightarrow\dfrac{\Lambda_{0}}{3}\beta^{a}\wedge\beta^{b}$. It is worthwhile to notice that since by construction $S_{G}$ is quadratic in the curvature, $S_{G}$ turns out to be invariant under the (more general) duality symmetry $\bar{\Omega}\leftrightarrow-\bar{\Omega}$ which splits as:
@@
In addition, a direct comparison with the two perturbative expansion parameters $\alpha$$\beta$ given in [19] yields:
@@
As for $k$, the case of a dynamical $\ell$ is studied in [12].
We can now consider $e=\dfrac{8\pi^{2}\alpha_{1}\ell^{2}}{k}=\dfrac{3\pi}{2\Lambda_{0}G}$ and $\ell^{2}=-\dfrac{3k}{\Lambda_{0}}$ as fixed. Upon identifying the B-I parameter as $\gamma=\dfrac{e}{r+2y}$, only two degrees of freedom corresponding to the relative value of the parameters $r$ and $y$ and to $|k|$ (the absolute value of $k$) are left.
@@
The Pontrjagin density and other terms described in (42) correspond to the Chern-Simons modified gravity [27] terms with a constant scalar field $\vartheta=-\big(\dfrac{2Gr}{\pi}+\dfrac{3}{\Lambda_{0}\gamma}\big)\dfrac{\alpha_{1}}{2}$ and potential term for $\vartheta$, $S_{\vartheta}=\int_{\mathscr{M}}V(\vartheta)$}, with:
@@
The Chern-Simons term also occurs in Loop Quantum Gravity (LQG) and String theory as mentionned in [27]. The fact that $\vartheta$ is a constant in our formalism leads to the speed of gravitational waves being the same as the speed of light as discussed in [28, 29] and is consistent with the gravitational wave GW170817 event.
@@
III Invariant polynomials and action for the Möbius group $G=SO(4,2)/\{\pm I\}$
@@
The generators of the above graded algebra are the same as the ones described previously in (II.1) except for $\mathfrak{g}_{0}\simeq\mathfrak{co}(3,1)$ which is generated by the $\{J_{ab}\}$’s and $\mathds{1}_{4}$. An element of $\mathfrak{co}(3,1)\simeq\mathfrak{g}_{0}$ is $\gamma_{\mathfrak{co}}=\frac{1}{2}\gamma_{\mathfrak{so}}^{ab}J_{ab}-z\mathds{1}_{4}$ with $\gamma_{\mathfrak{so}}\in\mathfrak{so}(3,,1)$ and $z\in\mathbb{R}$.
III.1 The Möbius Lie algebra as a symmetric Lie algebra
[…]
@@ III.1 The Möbius Lie algebra as a symmetric Lie algebra
of the Möbius group is reductive with respect to $H$ and is also a symmetric Lie algebra $\mathfrak{g}=\mathfrak{co}(3,1)\oplus\mathfrak{m}$ where $\mathfrak{m}=\mathfrak{g}_{1}\oplus\mathfrak{g}_{-1}\simeq{\mathbb{R}^{3,1}}^{*}\oplus\mathbb{R}^{3,1}$. The commutation relations (see (LABEL:eq:gradedLa) and [15, Example 5.2 and Chapter XI]) are given by
$[\mathfrak{g}_{0},\mathfrak{g}_{0}]\subset\mathfrak{g}_{0},\qquad[\mathfrak{g}_{0},\mathfrak{m}]\subset\mathfrak{m},\qquad[\mathfrak{m},\mathfrak{m}]\subset\mathfrak{g}_{0}\$(57)
@@
for which we adopt the matrix presentation [^14] Accordingly, the Killing form on $\mathfrak{g}=\mathfrak{h}\oplus\mathfrak{m}$ is computed to be $K_{\mathfrak{g}}(\varphi,\varphi^{\prime})=\operatorname{Tr}(\operatorname{ad}_{\varphi}\circ\operatorname{ad}_{\varphi^{\prime}})=n\operatorname{Tr}(cc^{\prime})+2n\big(zz^{\prime}+\eta(b,a^{\prime})+\eta(a,b^{\prime})\big)$ where $n$ corresponds to the one coming from $\dim\mathfrak{so}(n-1,1)=n(n-1)/2$.
@@
Let us now consider the mutation [^15] See Definition II.1 given above. $\mathfrak{g}^{\prime}=\mu(\mathfrak{g})=\mu(\mathfrak{h}\oplus\mathfrak{m})=\mathfrak{h}\oplus\mu(\mathfrak{m})=\mathfrak{h}\oplus\mathfrak{m}^{\prime}$ given by:
@@
Let us choose $\gamma_{1}=\bar{\gamma}_{1}=\dfrac{1}{\ell}$, $\gamma_{2}=\bar{\gamma}_{2}=\dfrac{1}{\ell^{\prime}}$ (where $\ell$ and $\ell^{\prime}$ are non-null real numbers) [^16] In that case, the Killing form on $\mathfrak{g}^{\prime}=\mathfrak{h}\oplus\mathfrak{m}^{\prime}$ can be easily computed by performing the scale transformations $b\mapsto\gamma_{1}b$ and $a\mapsto\gamma_{2}a$ into the expression for $K_{\mathfrak{g}}$, so that $K_{\mathfrak{g}^{\prime}}(\varphi,\varphi^{\prime})=n\operatorname{Tr}(cc^{\prime})+2n\Bigl(zz^{\prime}+\gamma_{1}\gamma_{2}\bigl(\eta(b,a^{\prime})+\eta(a,b^{\prime})\bigr)\Bigr).$ .
@@ A Cartan connection $\varpi$ with this $G/H$ Cartan geometry can accordingly be decomposed as:
The geometry modeled on the present $8$-dimensional quotient $G/H$ describes in general an $8$-dimensional manifold $\widetilde{\mathscr{M}}$ invariant under Lorentz transformations (elements in $SO(3,1)$) and dilations (elements of the multiplicative group $\mathbb{R}^{+}$) with soldering form $\varpi_{\mathfrak{m}}$. In the next section, we will see how we can obtain a description of a $4$-dimensional submanifold $\mathscr{M}$ of $\widetilde{\mathscr{M}}$ by imposing a condition on the soldering form (the $\mathfrak{m}$-valued part of the Cartan connection).
@@
(see e.g. [30]). For further use, it is worthwhile to notice that the $\mathfrak{m},\mathfrak{m}]\subset\mathfrak{h}$ contribution to $\bar{\Omega}_{\mathfrak{h}}$ is given by
@@
If we choose a Cartan geometry such that $\Theta=0$, $f=0$ and Ric(F)=0(F)=0, [11, Chapter 7, Def 2.7] then, as described in [31, p.52], $\bar{\Pi}$ and $\bar{\alpha}$ become, respectively, the Cotton and Schouten tensors, while $F$ would correspond to the Weyl tensor. These tensors constitute the building blocks of conformal gravity.
III.2 Characteristic poylynomials for $G=SO(4,2)/\{\pm I\}$ and associated action
@@
where $\mathscr{M}_{i}$ is either the $8$-dimensional manifold $\widetilde{\mathscr{M}}$ defined by the Cartan geometry $G/H$ or a $4$-dimensional submanifold $\mathscr{M}\subset\widetilde{\mathscr{M}}$ defined by the distribution:
Proposition III.1.
@@
Hence, by the Frobenius theorem [11, Prop.5.3, p.81] the corresponding distribution defines $\mathscr{M}$ as an integrable submanifold of $\widetilde{\mathscr{M}}$. Therefore, there is a subclass of conformal Cartan connections characterizing $\mathscr{M}$; this subclass can be reached through this kind of "gauge fixing". [^17] We are very indebted to Thierry Masson for his key input in the construction of this integrable distribution defined by this specific Cartan connection. This ought to resemble a gauge fixing which amounts to freezing the degrees of freedom of the $\mathfrak{g}_{1}$-part to be those of the $\mathfrak{g}_{-1}$-sector. It can be linked with [11, Proof of Proposition 3.1, p.285] where the two tetrads are taken to be linearly dependent each other. To some extent, this selects a mutated Weyl geometry within the mutated Möbius one. One thus gets $\dim\mathscr{M}=11-7=4$. One may also consider manifolds of other dimensions by imposing other constraints that reduce the number of degrees of freedom of the soldering form $\varpi_{\mathfrak{m}}$.
@@
the action being invariant under each of these transformations. Particular transformations satisfying 67 are:
@@
While for $k\neq 0$, a direct comparison between $S^{\prime}$ (31) and $S_{G}$ (III.2) amounts to identifying the 6 coupling constants present in $S^{\prime}$ via the system:
@@
  • Action (III.2) gives back the equations of motion (36)-(37).
  • Adding the term $\dfrac{vk}{2\pi^{2}\ell\ell^{\prime}}\eta_{ab}T^{a}\wedge\lambda\wedge\beta^{b}$ in the action (III.2) leads to a second source of curvature expressed in terms of the spin density of matter, torsion and their variations for the Einstein equations.
III.2.1 Fermionic matter field action
@@
If one considers the usual covariant derivative $D_{\mathfrak{so}}=(\partial_{\mu}+\dfrac{1}{4}A_{ab,\mu}\gamma^{a}\gamma^{b})dx^{\mu}$, the matter action similar to (30) is then given by  [^20] Where we set $\kappa^{2}\zeta=\tfrac{(\ell\ell^{\prime})^{2}}{384k^{2}}$ such that the overall factor is arranged to stick with the Dirac action given in [18, p.197].:
@@
Let $C_{i}$ be the $i$-dimensional Clifford algebra. We can extend the isomorphism 29 between $\mathfrak{so}(n-1,1)$ and $\mathfrak{spin}(n-1,1)$ to the morphism $\Phi_{\varepsilon}:\mathfrak{gl}(n,\mathbb{R})\rightarrow C_{2}\oplus C_{0}$. Let $M=M^{rs}e_{r}\bar{e}_{s}\in\mathfrak{gl(n,\mathbb{R)}}$. We define:
@@ III.2.2 The total action
In the other case, using the Hodge star operator lets us show that $T^{a}\wedge\beta_{a}=0$ is equivalent to requiring the axial vector part of torsion to vanish ($\tilde{t}^{a}=\dfrac{1}{6}T^{b}_{~\mu\nu}\varepsilon^{a~\mu\nu}_{~b}=0$) as is described in [32, Table 3]. The end result being that the new torsion $T^{a}_{~bc}=Z^{a}_{~bc}+\delta^{a}_{[b}t^{\vphantom{a}}_{c]}$ can be expressed solely in terms of $20$ independent components instead of the usual $24$. Here $Z^{a}_{~bc}$ is a $3$-indices tensor and $t_{c}$ is a $4$-dimensional vector with respectively $16$ and $4$ independent components.
@@
Using the relation $\gamma=\dfrac{2\alpha_{1}}{2\alpha_{2}-\alpha_{5}}$, given in [17, p13], we identify $\gamma=\dfrac{e}{r+2y}$, with $\gamma$ the Barbero-Immirzi parameter.
@@
$G_{kc}+(\dfrac{1}{2\gamma}-\dfrac{\Lambda_{0}Gr}{3\pi})R^{ab}_{~~rs}\varepsilon^{rs}_{~~kb}\eta_{ac}+\Lambda_{0}\eta_{kc}=-\dfrac{1}{2\alpha_{1}}(\tau_{ck}-\dfrac{1}{24}\partial_{\mu}\mathfrak{s}^{ab}_{~~ab\nu}\varepsilon^{\mu\nu}_{~~\;ck}).$ (90)
@@
In this particular case the fact that Torsion is null implies that the Einstein tensor $G_{kc}$ is symmetric (see [33, p3] for more details). When taking into account that $\eta_{kc}$ is symmetric as well, the anti-symmetricity of $\partial_{\mu}\mathfrak{s}^{ab}_{~~ab\nu}\varepsilon^{\mu\nu}_{~~\;ck}$ implies that the energy-momentum tensor $\tau_{ck}$ sports an anti-symmetric part if $\partial_{\mu}\mathfrak{s}^{ab}_{~~ab\nu}\varepsilon^{\mu\nu}_{~~\;ck}\neq 0$. This confirms that we can consider a matter Lagrangian that depends on the spin connection $A$ as we explicitly did.
Separating the energy-momentum tensor $\tau_{ck}$ as $\tau_{ck}=\tau_{M,ck}-\rho_{\text{vac}}\eta_{ck}$ with $\rho_{\text{vac}}\eta_{ck}$ the part related to the vacuum energy density $\rho_{\text{vac}}\geq 0$ then leads to:
$G_{kc}+(\dfrac{1}{2\gamma}-\dfrac{\Lambda_{0}Gr}{3\pi})R^{ab}_{~~rs}\varepsilon^{rs}_{~~kb}\eta_{ac}+\Lambda\eta_{kc}=-\dfrac{1}{2\alpha_{1}}(\tau_{M,ck}-\dfrac{1}{24}\partial_{\mu}\mathfrak{s}^{ab}_{~~ab\nu}\varepsilon^{\mu\nu}_{~~\;ck})$ (91)
@@
$G_{kc}$ $+(\dfrac{1}{2\gamma}-\dfrac{\Lambda_{0}Gr}{3\pi})R^{ab}_{~~rs}\varepsilon^{rs}_{~~kb}\eta_{ac}+\Lambda_{0}\eta_{kc}$
$=-\dfrac{1}{2\alpha_{1}}\Bigl(\tau_{ck}-\dfrac{1}{24}(\eta_{cs}\beta^{s}_{\nu}\partial_{\mu}-T_{c,\mu\nu})\bigl(\mathfrak{s}^{ab}_{~~abr}+(\dfrac{1}{2\pi G\gamma}-\dfrac{\Lambda_{0}r}{3\pi^{2}})T^{m}_{~rm}\bigr)\varepsilon^{\mu\nu r}_{~~~\,k}\Bigr)$ (92)
@@
The secondary source of curvature $\dfrac{1}{48\alpha_{1}}(\eta_{cs}\beta^{s}_{\nu}\partial_{\mu}-T_{c,\mu\nu})\bigl(\mathfrak{s}^{ab}_{~~abr}+(\dfrac{1}{2\pi G\gamma}-\dfrac{\Lambda_{0}r}{3\pi^{2}})T^{m}_{~rm}\bigr)\varepsilon^{\mu\nu r}_{~~~\,k}$ in (92) is highly dependent on the value of $\Lambda_{0}$ and $r$. One can for example tune $r=\dfrac{3\pi}{2\Lambda_{0}G\gamma}$ such that it is now dominated by the term $\dfrac{1}{48\alpha_{1}}(\eta_{cs}\beta^{s}_{\nu}\partial_{\mu}-T_{c,\mu\nu})\mathfrak{s}^{ab}_{~~abr}\varepsilon^{\mu\nu r}_{~~~\,k}$.
@@
Replacing $e=\dfrac{8\pi^{2}\alpha_{1}\ell\ell^{\prime}}{k}=\dfrac{3\pi}{2\Lambda_{0}G}$; $\ell\ell^{\prime}=-\dfrac{3k}{\Lambda_{0}}$ and $y=\dfrac{3\pi}{4\Lambda_{0}G\gamma}-\dfrac{r}{2}$ in the action (III.2), we obtain:
@@
[…]